OpenStudy (anonymous):

If the sum of the measures of the interior angles of a convex polygon is 2160°, how many sides does the polygon have?

jimthompson5910 (jim_thompson5910):

S = 180(n - 2) 2160 = 180(n-2) now solve for n

OpenStudy (aivantettet26):



OpenStudy (aivantettet26):